Tom found the sum of 9 and 4 by using ten.
2nd grade math riddle HELP!!!!

Answer by ewatrrr(24785)

 
Hi,
If I am understanding correctly
9 + 4 = 9 + 1 + 3 = 10 + 3 = 13